OpenStudy (anonymous):

[Calculus 2] Solve the integral (question in comments).

OpenStudy (anonymous):

If someone could explain how to solve this for me, it would be great! :)

OpenStudy (amistre64):

that looks like a doozie

OpenStudy (amistre64):

hmm\[\int2^{-(x^{-1})}x^{-2}~dx\]

OpenStudy (amistre64):

might be a by parts using the base as your downer and the x^-2 as the upper

OpenStudy (amistre64):

what is the derivative of 2^-(1/x) ?

OpenStudy (anonymous):

It's not integration by parts- this is from test 1. We did by parts on test 2.

OpenStudy (amistre64):

in general \[Dx~[a^x]=a^xlna\] this might be an ugly looking ...reverse product rule

OpenStudy (amistre64):

2^-(1/x) = y ln(2^-(1/x)) = ln(y) -1/x ln(2) = ln(y) ln(2)/x^2 = y'/y ln(2) y/x^2 = y' ln(2) 2^-(1/x)/x^2 = y' all we are missing is a useful ln2

OpenStudy (amistre64):

\[\frac1{ln2}\int ln(2)\frac{2^{-1/x}}{x^2}dx=\frac1{ln2}~2^{-1/x}\]

OpenStudy (amistre64):

+C

OpenStudy (amistre64):

i spose youd refer to it as a u-sub?

OpenStudy (anonymous):

^Yeah, it looks more like that than anything else.

OpenStudy (anonymous):

Thanks for your help! :)